OpenStudy (anonymous):

For a wedding reception, Violet rented a tent and twelve tables. The tent cost forty dollars to rent. Altogether, Violet spent two hundred twenty dollars. What is the cost of renting one table?

OpenStudy (jdoe0001):

\(\large \begin{array}{rllll} tent\\ 12tables \\\hline\\ 220 \end{array}\quad tent=40\to \begin{array}{rllll} 40\\ 12tables \\\hline\\ 220 \end{array}\implies 40+12tables=220\) solve for "tables"